# LAKE-RED: Camouflaged Images Generation by Latent Background Knowledge Retrieval-Augmented Diffusion

### v2

@@ 3.2.1 Background Knowledge Retrieval
$\mathbf{h}_{i}$ $=a_{i}\mathbf{x}^{f}\mathbf{W}_{i}^{V},$
@@ 4.2 Comparison with the State-of-the-art Methods
[…]
Image Blending
AB  [10]${}_{03}$
$\mathcal{F}+\mathcal{B}$
117.11
0.0645
126.78
0.0614
133.89
0.0645
120.21
0.0623
CI [7]${}_{10}$
$\mathcal{F}+\mathcal{B}$
124.49
0.0662
136.30
0.7380
137.19
0.0713
128.51
0.0693
AdaIN  [21]${}_{17}$
$\mathcal{F}+\mathcal{B}$
125.16
0.0721
133.20
0.0702
136.93
0.0714
126.94
0.0703
DCI [53]${}_{20}$
$\mathcal{F}+\mathcal{B}$
130.21
0.0689
134.92
0.0665
137.99
0.0690
130.52
0.0673
LCGNet [29]${}_{22}$
$\mathcal{F}+\mathcal{B}$
129.80
0.0504
136.24
0.0597
132.64
0.0548
129.88
0.0550
Image Inpainting
TFill  [64]${}_{22}$
$\mathcal{F}$
63.74
0.0336
96.91
0.0453
122.44
0.0747
80.39
0.0438
LDM  [41]${}_{22}$
$\mathcal{F}$
58.65
0.0380
107.38
0.0524
129.04
0.0748
84.48
0.0488
RePaint-L [36]${}_{22}$
$\mathcal{F}$
76.80
0.0459
114.96
0.0497
136.18
0.0686
96.14
0.0498
[…]
$\mathcal{F}$
39.55 0.0212 88.70 0.0428 102.67
0.0625
64.27 0.0355
@@
Module Prams(M)normal-$\downarrow$ MAC(G)normal-$\downarrow$ FPS(Hz)normal-$\uparrow$ Overall
BKRM
RCEM
LMP

### v4

@@ 3.2.1 Background Knowledge Retrieval
$\mathbf{h}_{i}$ $=a_{i}\mathbf{E}_{g}\mathbf{W}_{i}^{V},$
@@ 4.2 Comparison with the State-of-the-art Methods
Image Blending AB  [10]03 $\mathcal{F}+\mathcal{B}$ 117.11 0.0645 126.78 0.0614 133.89 0.0645 120.21 0.0623
CI [7]10 $\mathcal{F}+\mathcal{B}$ 124.49 0.0662 136.30 0.7380 137.19 0.0713 128.51 0.0693
AdaIN  [21]17 $\mathcal{F}+\mathcal{B}$ 125.16 0.0721 133.20 0.0702 136.93 0.0714 126.94 0.0703
DCI [53]20 $\mathcal{F}+\mathcal{B}$ 130.21 0.0689 134.92 0.0665 137.99 0.0690 130.52 0.0673
LCGNet [29]22 $\mathcal{F}+\mathcal{B}$ 129.80 0.0504 136.24 0.0597 132.64 0.0548 129.88 0.0550
Image Inpainting TFill  [64]22 $\mathcal{F}$ 63.74 0.0336 96.91 0.0453 122.44 0.0747 80.39 0.0438
LDM  [41]22 $\mathcal{F}$ 58.65 0.0380 107.38 0.0524 129.04 0.0748 84.48 0.0488
RePaint-L [36]22 $\mathcal{F}$ 76.80 0.0459 114.96 0.0497 136.18 0.0686 96.14 0.0498
Ours23 $\mathcal{F}$ 39.55 0.0212 88.70 0.0428 102.67 0.0625 64.27 0.0355
@@
Module Prams(M)$\downarrow$ MAC(G)$\downarrow$ FPS(Hz)$\uparrow$ Overall